OpenStudy (anonymous):

Can anyone find out what x and y is to this problem? Using the method: Substitution.? 2.5y + 3y =21 Thanks

OpenStudy (texaschic101):

what ? There is not even an x in the equation. Please check your question again and see if you wrote it wrong.

OpenStudy (anonymous):

I ment to put down 3x not y so it would be 2.5y + 3x = 21

OpenStudy (anonymous):

Does that still not make sense? :o

OpenStudy (texaschic101):

2.5y + 3x = 21 3x = 21 - 2.5y x = (21 - 2.5y)/3 now sub (21 - 2.5y)/3 in for x 2.5y + 3(21 - 2.5y)/3 = 21 2.5y + 21 - 2.5y = 21 0 = 0 so x = 0 now sub 0 in for x 2.5y + 3x = 21 2.5y + 3(0) = 21 2.5y = 21 y = 8.4 check.... 2.5(8.4) + 3(0) = 21 21 = 21 (correct) answer : x = 0, y = 8.4 I think I did that right.....?

OpenStudy (anonymous):

Thanks!!!!(:

OpenStudy (texaschic101):

your welcome :)